### v1

# Limit cycles and invariant algebraic curves

Abstract.
We give lower bounds in terms of $n,$ for the number of limit cycles of polynomial vector fields of degree $n,$ having any prescribed invariant algebraic curve. By applying them when the ovals of this curve are also algebraic limit cycles we obtain a new recurrent property for the Hilbert numbers. Finally, we apply our results to two important families of models: Kolmogorov systems and a general family of systems appearing in Game Theory.
Key words and phrases:
Hilbert 16th problem; limit cycles; invariant algebraic curves; Kolmogorov systems; Game Theory models
2020 Mathematics Subject Classification:
Primary: 34C07. Secondary: 34A05; 37G15
1. Introduction and statement of the main results
Let $X=(P,Q)$ be a planar polynomial vector field and $\gamma$ be one of its orbits. We say that $\gamma$ is algebraic if it is contained in an algebraic set. That is, if there is a polynomial $C\in\mathbb{R}[x,y]$ such that $\gamma\subset C^{-1}(\{0\})$. For simplicity, we denote the partial derivatives of a polynomial $C$ by $C_{x}$ and $C_{y}$. If we choose $C=0$ to be irreducible, then this condition reads as
(1) $C_{x}P+C_{y}Q=KC.$
where is $K\in\mathbb{R}[x,y]$ is known as the cofactor of $C$. It is easy to see that if the degree of $X$ is $n$, then the degree of $K$ is at most n1n-1.
In general, given $C\in\mathbb{R}[x,y],$ not necessarily irreducible, we say that $C=0$ is an invariant algebraic curve for $X$ if $C$ satisfies equation (1) for some $K\in\mathbb{R}[x,y]$ of degree at most n1.n-1. The term invariant follows from the fact that if $C$ satisfies (1) and the set $C^{-1}(\{0\})$ is not empty, then it is invariant by the flow of $X$.
In particular, if $\gamma$ is a limit cycle (i.e. isolated periodic orbit) then it is said to be an algebraic limit cycle. If the curve $C=0$ containing $\gamma$ is irreducible and has degree $c>0$, then we say that $\gamma$ has degree $c$.
For an account of known results on planar polynomial vector fields with algebraic invariant curves see [7, Chapter $8$] and its references. The interest on this subject goes back to the seminal works of Darboux and Poincaré.
In this paper we study how many limit cycles the family of polynomial vector fields, with a prescribed invariant algebraic, can have. We obtain lower bounds for this maximum number of limit cycles depending on the Hilbert numbers and their known lower bounds. As a consequence, we also give a new recurrent property for the Hilbert numbers themselves. Before stating our results more precisely we recall several known properties of the Hilbert numbers.
Let $\mathcal{X}^{n}$ be the set of planar polynomial vector fields $X=(P,Q)$ of degree $n,$ i.e. $\max\{\deg P,\deg Q\}=n.$ Given $X\in\mathcal{X}^{n}$, let $\pi(X)\in\mathbb{Z}_{\geqslant 0}\cup\{\infty\}$ denote its number of limit cycles. Given $n\in\mathbb{N}$ the Hilbert number $\mathcal{H}(n)\in\mathbb{Z}_{\geqslant 0}\cup\{\infty\}$ is
${H}(n)=\sup\{\pi(X)\colon X\in\mathcal{X}^{n}\}.$
It has been named after David Hilbert due to his famous list of open problems for the $20th$ century [2], being the second part of the $16th$ problem about upper bounds for $\mathcal{H}(n)$. Nowadays, it is not even known if $\mathcal{H}(2)$ is finite. For small values of $n$, the known best lower bounds are given in Table 1.
Table 1. Summary of the known lower bounds of the Hilbert numbers.
$n$ 22 33 44 55 66 77 88 99 $10$ $n$
$\mathcal{H}(n)$ 44 $13$ $28$ $37$ $53$ $74$ $96$ $120$ $142$ $O(n^{2}\ln n)$
The value of $\mathcal{H}(n)$ for $n=2$ is obtained in [4, 19], for $n=3$ in [16], and for $n\geqslant 4$ in [18]. The asymptotic bound follows from the results of [6, 12, 15].
In the recent work [9], the authors proved that the Hilbert numbers are realizable by structurally stable vector fields and, as a byproduct, by hyperbolic limit cycles.
Let us state our results. Given $C\in\mathbb{R}[x,y]$, a point $p\in\mathbb{R}^{2}$ is a singular point of $C=0$ if $C(p)=0$ and $\nabla C(p)=0$. We say that $C$ is non-degenerated if $C=0$ has at most a finite number of singular points. If $C=0$ does not have singular points, it is non-singular. An oval of $C=0$ is a bounded connected component of this set without singular points of $C=0.$ Let $\mathcal{O}(C)\in\mathbb{Z}_{\geqslant 0}$ denote the number of ovals of $C=0$.
Let $\mathcal{X}_{C}^{n}\subset\mathcal{X}^{n}$ be the set of planar polynomial vector fields of degree $n$ having $C=0$ as an invariant algebraic curve. For $n\in\mathbb{N}$ we define the Hilbert-C number as
$\mathcal{H}_{C}(n)=\sup\{\pi(X)\colon X\in\mathcal{X}_{C}^{n}\}.$
We say that $X\in\mathcal{X}_{C}^{n}$ is non-degenerated if it has at most finitely many singularities at $C=0$. Let $\Sigma_{C}^{n}\subset\mathcal{X}_{C}^{n}$ denote the family of non-degenerated vector fields.
In our main result we use the Hilbert numbers $\mathcal{H}(n)$ to provide a lower bound for $\mathcal{H}_{C}$ in such way that every oval of $C=0$ is a hyperbolic limit cycle.
Theorem A.
Given $C\in\mathbb{R}[x,y]$ non-degenerated of degree $c>0$ and $n\in\mathbb{Z}_{\geqslant 0}$, the following statements hold.
  1. (a)
    If $\mathcal{H}(n)<\infty$, then there is $X\in\Sigma_{C}^{n+c}$ such that every oval of $C=0$ is a limit cycle and $\pi(X)\geqslant\mathcal{H}(n)+\mathcal{O}(C),$
  2. (b)
    If $\mathcal{H}(n)=\infty$, then for each $k\in\mathbb{N}$ there is $X_{k}\in\Sigma_{C}^{n+c}$ such that every oval of $C=0$ is a limit cycle and $\pi(X_{k})\geqslant k+\mathcal{O}(C),$
being all limit cycles of $X$ or $X_{k}$ hyperbolic. Hence,
$\mathcal{H}_{C}(n+c)\geqslant\mathcal{H}(n)+\mathcal{O}(C),$
and as a consequence, for $n$ big enough, $\mathcal{H}_{C}(n)$ increases at least as fast as $O(n^{2}\ln n)$.
We remark that if in Theorem A we look simply for $X\in\mathcal{X}_{C}^{n+c}$ satisfying $\pi(X)\geqslant\mathcal{H}(n)$, then the proof is much easier. Consider for instance the case $\mathcal{H}(n)<\infty.$ Take $Y\in\mathcal{X}^{n}$ such that $\pi(Y)=\mathcal{H}(n)$ and such that none of its limit cycles intersects the curve $C=0.$ If we define $X=CY\in\mathcal{X}_{C}^{n+c},$ then $\pi(X)=\pi(Y)$ but if $C=0$ is not empty, $X\notin\Sigma_{C}^{n+c},$ because the curve $C=0$ is full of singularities of $X.$
Next four results are corollaries of Theorem A and previous results on literature. The first two provide recurrent properties for the Hilbert type numbers.
Corollary 1.
Let $L_{n}$ be any lower bound of $\mathcal{H}(n)$. Then $\mathcal{H}_{C}(n+c)\geqslant L_{n}+\mathcal{O}(C)$.
From the above result, by taking the values of Table 1 we obtain
$\mathcal{H}_{C}(2+c)\geqslant 4+\mathcal{O}(C),\,\mathcal{H}_{C}(3+c)\geqslant 13+\mathcal{O}(C),\,\mathcal{H}_{C}(4+c)\geqslant 28+\mathcal{O}(C),\ldots$
Corollary 2.
Given $n\in\mathbb{N}$ and $m\geqslant 1$ it holds
(2) $\mathcal{H}(n+m)\geqslant\mathcal{H}(n)+\operatorname{Har}(m),$
where $\operatorname{Har}(m)=(m-1)(m-2)/2+\bigl[1+(-1)^{m}\bigr]/2.$ Moreover the vector field of degree $n+m$ constructed to prove (2) has $\mathcal{H}(n)+\operatorname{Har}(m)$ hyperbolic limit cycles and at least $\operatorname{Har}(m)$ of them are algebraic and of degree $m.$
We observe that (2) can be seen as a generalization of [9, Theorem $1$], where it was proved that $\mathcal{H}(n+1)\geqslant\mathcal{H}(n)+1$. Nevertheless notice that this last inequality is stronger than (2) when $m=1,$ as $\operatorname{Har}(1)=0.$
Next two corollaries give lower bounds for two very interesting families of models that have particular invariant algebraic curves.
The first deals with Kolmogorov systems. Such systems extend the classical Lotka–Volterra predator–prey systems [13], and write as
(3) $\dot{x}=xp(x,y),\quad\dot{y}=yq(x,y),$
with $p$, $q\in\mathbb{R}[x,y].$ Notice that for them, if we consider $C(x,y)=xy,$ then $C=0$ is a reducible invariant algebraic curve. In this case we write $\mathcal{H}_{C}$ as $\mathcal{H}_{K}$.
Corollary 3.
Let $\mathcal{H}_{K}(n)$ denote the maximum number of limit cycles of (3) in the first quadrant, where $p$ and $q$ vary among all polynomials of degree at most n1n-1. Next table gives their exact values for $n=1,2$ and some lower bounds for $n\geqslant 3$.
$n$ 11 22 33 44 55 66 77 88 $n$
$\mathcal{H}_{K}(n)$ 0 0 66 $13$ $22$ $28$ $37$ $53$ $O(n^{2}\ln n)$
Case $n=1$ is trivial, as linear systems do not have limit cycles. When $2\leqslant n\leqslant 5$, it follows from previous works, see the references in the proof. In particular, for $3\leqslant n\leqslant 5$ the values have been obtained by high order Andronov–Hopf bifurcations. The results for bigger $n$ are new and follow from our approach based on Theorem A.
By using the same tools we obtain a similar result for some models in Game Theory. For our goals it is sufficient to observe that when the game has two players it can modeled by vector fields of the form
(4) $\dot{x}=x(x-1)p(x,y),\quad\dot{y}=y(y-1)q(x,y),$
with $p$ and $q\in\mathbb{R}[x,y].$ For details about their construction, and the interpretation of $x$ and $y$ as strategies, see [8, 11] and their references. Clearly, these models have the reducible invariant algebraic curve C(x,y)=x(x1)y(y1)=0C(x,y)=x(x-1)y(y-1)=0 of degree $c=4.$ Moreover, the region of interest for applications is the open square $\mathcal{Q},$ delimited by the invariant four straight lines, which is also clearly invariant by the flow. To recall the shape of this region, in this case we write $\mathcal{H}_{C}=\mathcal{H}_{{\scriptscriptstyle{\square}}}$.
Corollary 4.
Let $\mathcal{H}_{\scriptscriptstyle{\square}}(n)$ denote the maximum number of limit cycles of (4) in the open square $\mathcal{Q}$, where $p$ and $q$ vary among all polynomials of degree at most n2n-2. Next table gives their exact values for $n=2,3$ and some lower bounds for $n\geqslant 4$.
$n$ 22 33 44 55 66 77 88 99 $10$ $n$
$\mathcal{H}_{\scriptscriptstyle{\square}}(n)$ 0 11 55 55 55 $13$ $28$ $37$ $53$ $O(n^{2}\ln n)$
As in Corollary 4, the proof for $n=2$ is easy: it is an integrable system. Cases $3\leqslant n\leqslant 4$ follow from previous results. Again the main tool when $n=4$ is the study of a high order Andronov–Hopf bifurcation. The results for $n=5$ and $6$ follow trivially from the one for $n=4,$ and for bigger $n$ they are new and a consequence of the main result of this work. It is clear that the cases $n=5$ and $6$ deserve more attention. It is sure that a careful study of them will give higher lower bounds.
The paper is organized as follows. In Section 2 we give some preliminary results about properties of the Hilbert numbers and about prescribed algebraic limit cycles. The main theorem and its corollaries are proved in Section 3.
2. Preliminaries
The first result that we will use is a simplified version of [9, Theorem $2$].
Theorem 1 ([9]).
For $n\in\mathbb{N}$, the following statements hold.
  1. (a)
    If $\mathcal{H}(n)<\infty$, then there is $X\in\mathcal{X}^{n}$ having $\mathcal{H}(n)$ hyperbolic limit cycles.
  2. (b)
    If $\mathcal{H}(n)=\infty$, then for each $k\in\mathbb{N}$ there is $X_{k}\in\mathcal{X}^{n}$ having at least $k$ hyperbolic limit cycles.
We will also need an adaptation of [5, Theorem 1]. Although its proof is essentially the same, for sake of self-completeness we include it here.
Theorem 2.
Let $C\in\mathbb{R}[x,y]$, be non-degenerated and of degree $c>0$, $D\in\mathbb{R}[x,y]$ a polynomial of degree one that does not divide $C$ and such that the $D=0$ does not intersect any oval of $C=0$. Let $\alpha$, $\beta\in\mathbb{R}$ such that $\alpha D_{x}+\beta D_{y}\neq 0$. Then the polynomial vector field $Y=(P,Q)$ given by
(5) $P=\alpha C-DC_{y},\quad Q=\beta C+DC_{x},$
has all the ovals of $C=0$ as hyperbolic limit cycles and $Y\in\Sigma_{C}^{c}$.
Proof.
It is clear that $C=0$ is an invariant algebraic curve for (5) because $C_{x}P+C_{y}Q=\big(\alpha C_{x}+\beta C_{y}\big)C.$ Moreover, clearly $Y\in\Sigma_{C}^{c}$. Let $\gamma$ be an oval of $C=0$, i.e. a bounded connected component of $C=0$ without singular points. Since $D=0$ does not intersect $\gamma$ and $\nabla C$ never vanishes at it, we have that $Y$ does not have singularities at $\gamma$ and thus it is a periodic orbit of $Y$. It can be parameterized by time and let $T$ denote its minimal period. To prove that it is a hyperbolic limit cycle it suffices to show that next integral does not vanishes, see [7, Section 7.3]. We have,
$\displaystyle\int_{0}^{T}\operatorname{div}(Y)\,{\rm d}t$ $=\displaystyle\int_{0}^{T}P_{x}+Q_{y}\,{\rm d}t=\int_{0}^{T}(\alpha+D_{y})C_{x}+(\beta-D_{x})C_{y}\;{\rm d}t\vskip 5.69046pt$
$\displaystyle\stackrel{{\scriptstyle(*)}}{{=}}\int_{0}^{T}\frac{1}{D}\bigl[(\alpha+D_{y})\underbrace{(\beta C+DC_{x})}_{{\rm d}y/{\rm d}t}-(\beta-D_{x})\underbrace{(\alpha C-DC_{y})}_{{\rm d}x/{\rm d}t}\bigr]{\rm d}t\vskip 5.69046pt$
$=\int_{\gamma}\frac{\alpha+D_{y}}{D}{\rm d}y-\frac{\beta-D_{x}}{D}{\rm d}x=\pm\int_{\Gamma}\frac{\alpha D_{x}+\beta D_{y}}{D^{2}}\;{\rm d}x{\rm d}y\neq 0.$
Notice that in the last equality we used Green’s theorem, $\Gamma$ denotes the interior region bounded by $\gamma,$ the sign in front of the last integral depends on the time orientation of $\gamma$ and in ()(*) we have used that in the integrated function we can take $C=0$ because $\gamma$ is contained in this set. ∎
3. Proofs
Proof of Theorem A.
Suppose first $\mathcal{H}(n)<\infty$. It follows from Theorem 1(a)(a) that there is $X=(U,V)$ such that $\pi(X)=\mathcal{H}(n)$ and all its limit cycles are hyperbolic. Let $B\subset\mathbb{R}^{2}$ be a closed ball containing all the limit cycles of $X$ in its interior. Translating and scaling $X$ if necessary, we can suppose $B\cap C^{-1}(\{0\})=\emptyset$.
It follows from Theorem 2 that there is $Y=(P,Q)\in\Sigma_{C}^{c}$ having every oval of $C$ as hyperbolic limit cycle. Recall that then $C_{x}P+C_{y}Q=KC,$ for some cofactor $K.$ Given $\varepsilon>0$ let $Z=(R,S)$ be given by
$R=CU+\varepsilon P,\quad S=CV+\varepsilon Q.$
Notice that $C_{x}R+C_{y}S=\big(C_{x}U+C_{y}V+\varepsilon K\big)C$ and thus $Z\in\mathcal{X}_{C}^{n+c}$. By Theorem 2, if $\gamma$ is one of the limit cycles of $Y$, given by one of the ovals of $C,$ it is hyperbolic. Therefore, for $\varepsilon>0$ small enough it is also a hyperbolic limit cycle for $Z.$ Since all the limit cycles of $X$ are inside $B,$ are also hyperbolic, and $B\cap C^{-1}(\{0\})=\emptyset$, it follows that they persist and give rise to nearby hyperbolic limit cycles for $Z$ when $\varepsilon>0$ is again small enough. Hence there is $\varepsilon>0$ small enough such all the algebraic limit cycle of $Y$ remain as limit cycles of $Z$ and the hyperbolic limit cycles of $X$ also persist as limit cycles of $Z$, being small pertubations of the original ones, proving
(6) $\mathcal{H}_{C}(n+c)\geqslant\pi(Z)\geqslant\pi(X)+\pi(Y)=\mathcal{H}(n)+\mathcal{O}(C).$
Observe now that $Z|_{C=0}=\varepsilon Y|_{C=0}$ and thus $Z\in\Sigma_{C}^{n+c}$.
If $\mathcal{H}(n)=\infty$, then we have from Theorem 1(b)(b) that for each $k\in\mathbb{N}$ there is $X_{k}\in\Sigma^{n}$ with at least $k$ hyperbolic limit cycles. Applying the previous reasoning on each $X_{k}$ we obtain a sequence $\{Z_{k}\}\subset\Sigma_{C}^{n+c}$, such that $\pi(Z_{k})\geqslant k+\mathcal{O}(C)$.
It follows from [12, Theorem $2.1$] that there is $M>0$ such that for each $n\geqslant 3$, $\mathcal{H}(n)\geqslant(n^{2}\log_{2}n)/2-Mn^{2}\log_{2}\log_{2}n.$ Hence from (6) we have that, for $n$ big enough, $\mathcal{H}_{C}(n)$ increases at least as fast as $O(n^{2}\ln n)$. ∎
Proof of Corollary 1.
From (6) we have $\mathcal{H}_{C}(n+c)\geqslant\mathcal{H}(n)+\mathcal{O}(C)\geqslant L_{n}+\mathcal{O}(C).$
Proof of Corollary 2.
For each $m\in\mathbb{N}$ we have from Harnack’s Curve Theorem [20] that there is $C\in\mathbb{R}[x,y],$ with $C=0$ irreducible and of degree $m>0,$ with $\operatorname{Har}(m)$ ovals. The proof follows from Theorem A applied to this curve. ∎
Proof of Corollary 3.
We have from [1] and [17] that $\mathcal{H}_{K}(2)=0$ and $\mathcal{H}_{K}(3)\geqslant 6$. From [3] we know that $\mathcal{H}_{K}(4)\geqslant 13$ and $\mathcal{H}_{K}(5)\geqslant 22$.
From Theorem A we have $\mathcal{H}_{K}(n+2)\geqslant\mathcal{H}(n)$ and observe that from the proof of this theorem we can choose these limit cycles to be at the first quadrant. Hence the corollary follows by using the known lower bounds of $\mathcal{H}(n)$ given in Table 1.
Finally we observe that the vector fields provided by Theorem A are non-degenerated, i.e. they have at most finitely many singularities at $xy=0$. ∎
Proof of Corollary 4.
I We know from [1] and [14, Theorem $2.5.4$] that $\mathcal{H}_{\scriptscriptstyle{\square}}(2)=0$ and $\mathcal{H}_{\scriptscriptstyle{\square}}(3)=1$. From [8, 10] we have $\mathcal{H}_{\scriptscriptstyle{\square}}(4)\geqslant 5$. Trivially, $\mathcal{H}_{\scriptscriptstyle{\square}}(5)$ and $\mathcal{H}_{\scriptscriptstyle{\square}}(6)$ are at least $5$. From Theorem A we know that $\mathcal{H}_{\scriptscriptstyle{\square}}(n+4)\geqslant\mathcal{H}(n)$ and, as in the previous corollary, the results for $n\geqslant 7$ follow from the known lower bounds for $\mathcal{H}(n)$.
Moreover, all these limit cycles are hyperbolic and can be located to be realizable by the game, that is, in the open square $\mathcal{Q}.$
Acknowledgments
This work is supported by the Spanish State Research Agency, through the project PID2022-136613NB-I00, and by São Paulo Research Foundation (FAPESP), Brazil, grants 2021/01799-9 and 2024/15612-6.